### v1

# NOCTA: Non-Greedy Objective Cost-Tradeoff Acquisition for Longitudinal Data

Abstract
In many critical applications, resource constraints limit the amount of information that can be gathered to make predictions. For example, in healthcare, patient data often spans diverse features ranging from lab tests to imaging studies. Each feature may carry different information and must be acquired at a respective cost of time, money, or risk to the patient. Moreover, temporal prediction tasks, where both instance features and labels evolve over time, introduce additional complexity in deciding when or what information is important. In this work, we propose NOCTA, a Non-Greedy Objective Cost-Tradeoff Acquisition method that sequentially acquires the most informative features at inference time while accounting for both temporal dynamics and acquisition cost. We first introduce a cohesive estimation target for our NOCTA setting, and then develop two complementary estimators: 1) a non-parametric method based on nearest neighbors to guide the acquisition (NOCTA-NP), and 2) a parametric method that directly predicts the utility of potential acquisitions (NOCTA-P). Experiments on synthetic and real-world medical datasets demonstrate that both NOCTA variants outperform existing baselines.
1 Introduction
In many real-world scenarios, acquiring features at inference time comes at a cost. For example, in clinical settings (a driving application for this work), diagnostic decisions often rely on sequentially gathering information such as lab tests or imaging. Each type of data acquisition comes at the cost of time, financial expense, or risk to the patient. In contrast, many of the existing machine learning frameworks assume that all features are freely available at the outset. In this work, we focus on a longitudinal Active Feature Acquisition (AFA) (Saar-Tsechansky et al., 2009; Kossen et al., 2022) setting, where a model is responsible for: (1) sequentially deciding what features to acquire and at what time during inference, balancing their utility against acquisition costs, and (2) making predictions with the partially observed features it has chosen to acquire.
In particular, we focus on a longitudinal setting, where long temporal contexts may be crucial for both selecting valuable feature acquisitions and making accurate predictions. Fig. 1 illustrates an example of longitudinal AFA methods driven by an autonomous agent operating in a clinical scenario. At each visit, the agent (1) reviews previously collected data alongside any measurement acquired during the current visit, (2) predicts the patient’s current status, and (3) recommends the next follow-up by specifying both the visit interval and the informative tests to perform, while accounting for acquisition costs. Although in principle a comprehensive set of tests could be performed, practical constraints such as time, cost, and resource limitations necessitate a more selective approach. As a result, the agent prioritizes the most essential and informative acquisitions. Between the visits, the agent projects the patient’s health trajectory to future time points, using the data acquired so far and the length of the proposed time interval. Because the agent cannot revisit earlier time points, it must carefully decide whether to defer or perform tests at each time step to avoid missing potentially important acquisitions.
Figure 1: Analogy to longitudinal active feature acquisition in a clinical scenario: at each time point, an autonomous agent reviews previously collected and newly acquired data, predicts the patient’s current status, and recommends a subset of examinations for the next visit. This process is repeated until no further follow-up is recommended by the agent.
This longitudinal AFA setting presents several challenges, including: deciding which features to acquire at each visit, facilitating early predictions for timely interventions, and accounting for temporal settings where missed acquisitions at earlier time points are permanently inaccessible. Although previous work has explored various approaches for longitudinal AFA, some methods (Kossen et al., 2022) make a single prediction across all time points, neglecting the need for dynamic predictions, which is critical in clinical settings to catch deterioration early and adjust the treatment plans. Furthermore, recent reinforcement learning (RL)-based methods (Qin et al., 2024; Nguyen et al., 2024) often face challenges in optimization, while greedy strategies (Ghosh and Lan, 2023; Gadgil et al., 2024) risk suboptimal decisions by failing to anticipate the joint information of potential future acquisitions.
Contributions. In this work, we propose NOCTA, Non-Greedy Objective Cost-Tradeoff Acquisition, a non-greedy method that explicitly addresses longitudinal AFA challenges without training an RL policy, while maintaining an effective acquisition strategy. Firstly, we introduce the estimation target of NOCTA that directly balances prediction accuracy against the cost of acquiring features. Secondly, we develop two alternative approaches within the NOCTA framework: (1) a non-parametric approach that leverages nearest neighbors to guide the acquisition (NOCTA-NP), and (2) a parametric approach that directly estimates the utility of future acquisitions (NOCTA-P). Thirdly, we evaluate our method on both synthetic and real-world medical datasets, demonstrating that NOCTA consistently outperforms state-of-the-art baselines while achieving lower overall acquisition cost.
2 Related Work
Active Feature Acquisition. Prior studies in AFA (Saar-Tsechansky et al., 2009; Sheng and Ling, 2006) have considered the trade-off between accurate prediction and the cost of acquiring features. Recent works (Shim et al., 2018; Yin et al., 2020; Janisch et al., 2020) frame AFA as an RL problem; however, they are challenging to train in practice due to the complicated state and action space. Other works (Li and Oliva, 2021) consider using surrogate generative models, which model complicated multidimensional distributions, or employ greedy policies that cannot account for the joint informativeness of features (Covert et al., 2023; Ma et al., 2018; Gong et al., 2019). To overcome these issues, (Valancius et al., 2023) employs a non-parametric oracle-based approach that balances the estimated prediction loss against acquisition costs. Alternatively, (Ghosh and Lan, 2023) learn a non-greedy, differentiable acquisition policy by jointly optimizing a selection network and a prediction model through end-to-end training.
Longitudinal Active Feature Acquisition. Longitudinal AFA is an extension of standard AFA that incorporates the temporal dimension. In this setting, a policy must account for temporal constraints where past time points can no longer be accessed. As a result, a policy must decide which features to acquire and determine the optimal timing for each acquisition. For instance, (Zois and Mitra, 2017) derives dynamic-programming and low-complexity myopic policies for finite-state Markov chains, ASAC (Yoon et al., 2019) uses an actor-critic approach to jointly train a selector network to select features and a predictor network for the prediction. Additionally, (Nguyen et al., 2024) uses reinforcement learning to optimize the timing of follow-up acquisitions, but acquires all available features at the selected time point. Recently, (Qin et al., 2024) prioritizes timely detection by allowing flexible follow-up intervals in continuous-time settings, ensuring predictions are timely and accurate. However, approaches that rely on reinforcement learning can be challenging to train due to issues such as sparse rewards, high-dimensional observation spaces, and complex temporal dependencies Li and Oliva (2021). Rather than relying on reinforcement learning to learn a policy, our approach explicitly models and optimizes the trade-off between predictive accuracy and acquisition cost.
3 Methodology
Figure 2: An overview of the NOCTA framework. Starting at time step $t$ with the observed features $x_{o}$ and a set of candidate acquisition masks, NOCTA provides two complementary estimators to predict the accuracy–cost tradeoff for each candidate. It then selects the mask with the minimum estimated tradeoff value, predicts for the time points in between, and acquires the corresponding features $x_{v}$, which are appended to $x_{o}$. This updated set of features is again passed to the predictor, and this process repeats at the next time step $t^{\prime}>t$.
Notation. We formally define the longitudinal AFA setup: let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ denote the dataset of $N$ instances (e.g., patients). For each $i$, $x_{i}$ is a temporal sequence with $L$ time steps, i.e., $x_{i}=\{x_{i,t}\}_{t=1}^{L}$, with $t\in\{1,\ldots,L\}$ indexing the time step. At each time step, $x_{i,t}$ includes measurements of $M$ features (note that it is possible to extend to individual acquisitions that yield multi-dimensional feature vectors for multi-modal cases), i.e., $x_{i,t}=\{x^{m}_{i,t}\}_{m=1}^{M}$, and $x^{m}_{i,t}$ contains measurements of the $m$-th feature. Note that it is possible for $x^{m}_{i,t}=\emptyset$, which indicates the feature is unobserved. In this work, we focus on a classification task where labels $y_{i,t}\in\{1,2,\ldots,C\}$ are defined for each timepoint $t$, and $C$ represents the total number of classes. Following (Qin et al., 2024; Yoon et al., 2019), we consider the cost of measuring the $m$-th feature to be time-invariant, denoted as $c^{m}$. Lastly, we use colons to represent selecting subsequences along a specific axis; for example, $x_{1:t}=(x_{1},\ldots,x_{t})$ represents the measurements from timepoint $1$ to $t$.
3.1 Longitudinal AFA modeled as an MDP
Longitudinal AFA can be formulated as a Markov Decision Process (MDP). The state $s_{t}=\bigl{(}{x}_{o},\,o,\,t\bigr{)}$, where ${x}_{o}$ denotes the (partial) observations before time $t$ and $o\subseteq\{(m,t^{\prime})\mid m\in\{1,\ldots,M\},\,t^{\prime}\in\{1,\ldots,t-1\}\}$ records which features were acquired at each past step. We suppose access to a pre-trained arbitrary conditioning classifier $\hat{y}$  (Shim et al., 2018; Li and Oliva, 2021) , that can predict the label at any timepoint $k$ with $x_{o}$: $\hat{y}_{k}(x_{o})$. The action space is $a\in\{a_{t^{\prime}}\}_{t^{\prime}=t}^{L}\cup\{\varnothing\}$, where $a_{t^{\prime}}\in\{0,1\}^{M}$ indicates which of the $M$ features to acquire at the current or future time $t^{\prime}$, or, when $a=\varnothing$, to terminate and predict for all remaining timepoints. For a non-terminal action, the transition is $\bigl{(}{x}_{o},\,o,\,t\bigr{)}\rightarrow\bigl{(}{x}_{o\cup a_{t^{\prime}}},\,o\cup a_{t^{\prime}},\,t^{\prime}+1\bigr{)}$, where (for notational simplicity) $o\cup a_{t^{\prime}}$ denotes the union of the previous acquisitions and tuples of features at time $t^{\prime}$ in $a_{t^{\prime}}$. Since this implies that we are not acquiring any information until time $t^{\prime}$, the reward is the negative cross-entropy (CE) loss of making predictions up to $t^{\prime}$ ($\hat{y}_{t},\ldots,\hat{y}_{t^{\prime}-1}$) with previous information (${x}_{o}$) and respective negative CE loss and cost of acquiring new features at time $t^{\prime}$: $-\sum_{k=t}^{t^{\prime}-1}\mathrm{CE}(\hat{y}_{k}({x}_{o}),\,y_{k}\bigr{)}-\mathrm{CE}(\hat{y}_{t^{\prime}}({x}_{o\cup a_{t^{\prime}}}),\,y_{t^{\prime}}\bigr{)}-\alpha\sum_{m=1}^{M}a_{t^{\prime},m}c^{m},$ where $\alpha$ is an accuracy-cost tradeoff hyperparameter. When terminated ($a=\varnothing$) at timepoint $t$, we are not acquiring any additional information for the remaining timepoints, thus the reward is $-\sum_{k=t}^{L}\mathrm{CE}(\hat{y}_{k}({x}_{o}),\,y_{k}\bigr{)}$, which represents the negative cross-entropy loss accumulated for all timepoints starting from $t$ to $L$. We denote our policy as $\pi\bigl{(}{x}_{o},\,o,\,t\bigr{)}$, which selects actions to maximize the expected cumulative rewards over timepoints through $L$.
Challenges. The challenges of longitudinal AFA with MDP include: 1) large and complex action space, which requires decisions on when to acquire and which features to acquire; 2) complicated state space, as it evolves dynamically as new features are captured over time; and 3) credit assignment problem, since acquisitions are awarded at an aggregate level, it is hard to disentangle exactly which of the acquisitions were most responsible for prediction rewards (Li and Oliva, 2021). Despite these, existing works (Qin et al., 2024; Yoon et al., 2019; Kossen et al., 2022; Nguyen et al., 2024) often rely on reinforcement learning. In this work, we propose NOCTA, an RL-free approach that is not only easier to optimize but also performs well in practice.
3.2 NOCTA: Generalized Cost Objective
Algorithm 1 NOCTA: Non-Greedy Objective Cost-Tradeoff Acquisition
1:Observed features $o$ (possibly $o=\varnothing$), instance values $x_{o}$, number of features $M$, number of time steps $L$, cost for feature $m$ as $c^{m}$, tradeoff parameter $\alpha>0$, estimator $\hat{y}_{k}$.
2:Initialize: predictions [$\leftarrow$[\,]← [ ]; terminate $\leftarrow$ false; $t\leftarrow 0$
3:while not terminate do
4:    $\mathcal{V}\leftarrow\{(m,t^{\prime})\mid m\in\{1,\ldots,M\},\,t^{\prime}\in\{t+1,\ldots,L\}\}\setminus o$
5:    $\hat{u}(x_{o},o)\approx\underset{v\subseteq\mathcal{V}}{\arg\!\min}\Bigl{[}\mathbb{E}_{y_{t+1:L},x_{v}\mid x_{o}}\,[\ell(x_{o}\cup x_{v},y,t)]+\alpha\sum_{m\in v}c^{m}\Bigr{]}$
6:    if $\hat{u}(x_{o},o)=\varnothing$ then
7:        $t_{\text{next}}\leftarrow L$; terminate $\leftarrow$ true
8:    else
9:        $t_{\text{next}}\leftarrow\min\{\,t^{\prime}\mid(m,t^{\prime})\in\hat{u}(x_{o},o)\}$
10:        $o\leftarrow o\cup\{(m,t_{\text{next}})\mid(m,t_{\text{next}})\in\hat{u}(x_{o},o)\}$
11:    end if
12:    for $t^{\prime}=t+1,\dots,t_{\text{next}}$ do
13:        append $\hat{y}_{t^{\prime}}\bigl{(}x_{o_{1:t^{\prime}}}\bigr{)}$ to predictions
14:    end for
15:    $t\leftarrow t_{\text{next}}$
16:end while
17:return predictions
We now introduce NOCTA through a novel objective that shall act as a proxy to the MDP value of states. Let $\mathcal{V}=\{(m,t^{\prime})\mid m\in\{1,\ldots,M\},\,t^{\prime}\in\{t+1,\ldots,L\}\}$ represent the set of candidate feature-time pairs that are still available to acquire after time $t$. We ask: (1) which subset in $\mathcal{V}$ should we acquire to balance the tradeoff between improved accuracy and acquisition cost, or (2) should we terminate and make the prediction for the remaining time points based on what we have observed in $x_{o}$? That is, NOCTA determines the set of potential future acquisitions $u$:
$u\bigl{(}{x}_{o},o\bigr{)}=\underset{v\subseteq\mathcal{V}}{\arg\!\min}\,\mathbb{E}_{y_{t+1:L},x_{v}\mid{x}_{o}}[\ell(x_{o}\cup x_{v},y,t)]+\alpha\!\!\!\!\sum_{(m,t^{\prime})\in v}\!\!\!\!c^{m},$ (1)
where
$\ell(x_{o}\cup x_{v},y,t)=\sum_{k=t+1}^{L}\mathrm{CE}\bigl{(}\hat{y}_{k}({x}_{o}\cup{x}_{v_{t+1:k}}),\,y_{k}\bigr{)},$ (2)
represents the accumulated loss when only acquiring the set of future feature-time pairs $v$, and ${x}_{v_{t+1:k}}$ denotes the features in the candidate subset $v$ from time step $t+1$ to time step $k$. Because the full trajectories $(x_{v},y_{t+1:L})$ are available for the training set $\mathcal{D}$, the expectation in Eq. (1) can be evaluated during training. Note that predictions can only rely on the current and previously captured data, but not on that which will be obtained in the future. Intuitively, Eq. (1) non-greedily looks ahead at how acquiring a new set of feature-time pairs in $\mathcal{V}$ might reduce the classification loss from $t+1$ to $L$, while penalizing the acquisition cost. For simplicity, we refer to feature-time pairs as "features" in the remainder of the paper.
When Eq. (1) returns $u\bigl{(}x_{o},o\bigr{)}=\varnothing$, it indicates that the improvement in the loss is not worth the cost of acquiring a new feature, and thus we should stop acquiring and predict for all subsequent time points using what has already been observed. Otherwise, we can acquire new $u\bigl{(}x_{o},o\bigr{)}=v\subseteq\mathcal{V}$, make the prediction for the current and subsequent time points until the next acquisitions occur. This non-greedy policy for NOCTA directly balances the tradeoff between acquisition cost and predictive performance. The acquisition process of NOCTA is further described in Algorithm 1. Note, however, that in practice we do not have access to the true expectation of labels and unacquired values, $\mathbb{E}_{y_{t+1:L},x_{v}\mid{x}_{o}}$. In this paper, we consider two ways to estimate $u\bigl{(}{x}_{o},o\bigr{)}$: a non-parametric method (Sec. 3.3) and a parametric method (Sec. 3.4).
Following Valancius et al. (2023), one may connect our NOCTA acquisition objective (Eq. (1)) to the full RL optimization via a lower bound. In other words, acquiring according to the NOCTA objective (Eq. (1)) serves as an approximation to the optimum longitudinal AFA MDP. Additional details can be found in Appendix A.4.
3.3 NOCTA-NP: Non-parametric Method
In this subsection, we show how to estimate $u\bigl{(}{x}_{o},o\bigr{)}$ using a non-parametric method, coined NOCTA-NP. The method uses the observed features and estimates Eq. (1) by using similar cases in the training set identified via nearest neighbors. Furthermore, in order to assuage issues with the curse of dimensionality, we propose to compute nearest-neighbor distances based on representations from an embedding network, which is trained to encode information for future candidate selection.
Estimating $u\bigl{(}{x}_{o},o\bigr{)}$. We estimate $u\bigl{(}{x}_{o},o\bigr{)}$ at the current time $t$ by sampling labels and unacquired features through neighbors:
$\mathbb{E}_{y_{t:L},x_{v}\mid{x}_{o}}\bigl{[}\ell\bigl{(}x_{o}\cup x_{v},y,t\bigr{)}\bigr{]}\approx\frac{1}{K}\sum_{i\in N_{K}(x_{o})}\sum_{k=t+1}^{L}\mathrm{CE}(\hat{y}_{k}({x}_{i,o}\cup{x}_{i,v_{1:k}}),\,y_{i,k}\bigr{)},$ (3)
where $N_{K}(x_{o})$ represents the set of $K$-nearest neighbors in the dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, restricted to only the observed features $o$. That is, Eq. (3) leverages the neighbors to approximate the expected cross-entropy loss for any candidate acquisition plan $v$. The neighbors could be determined by comparing the distance $d(x_{o},x_{i,o})\to\mathbb{R}_{+}$, which directly uses the raw feature values, for example $d(x_{o},x_{i,o})=\|x_{o}-x_{i,o}\|_{2}$. However, note that $x_{o}$ accrues features through multiple time points, leading to a high-dimensional space where neighbors may not be informative. Hence, we propose computing distance using embedding network representations.
Given access to any future observations $x_{v}$ of the training set $\mathcal{D}$, we train an embedding network representation so that: for any two instances sharing the same set of observed features $o$ are close whenever additional acquisitions affect their accuracy-cost tradeoff objective in a similar way. Specifically, for each training sample $(x_{i},y_{i})\in\mathcal{D}$ with observed features $o$ (e.g., $x_{i,o}$), we define a ‘future-candidate’ distribution over candidate acquisition subsets, weighting each subset by its accuracy-cost tradeoff. The embedding network is then trained via a contrastive-type loss (Chopra et al., 2005) that pulls together instances whose future-candidate distributions are similar (i.e., whose candidate acquisitions have similar accuracy-cost tradeoff impacts). We formally define the future-candidate distribution and embedding network below:
Future-Candidate Distribution $p(\cdot\mid x_{o})$. For a partially observed features $x_{o}$ at time $t$ and corresponding labels $y$ of the training dataset $\mathcal{D}$, we collect its top-$\kappa$ candidate subsets ${v_{1},\ldots,v_{\kappa}}\subseteq\mathcal{V}$, in ascending order of:
$r_{l}=\ell(x_{o}\cup x_{v_{l}},y,t)+\alpha\!\!\!\!\sum_{(m,t^{\prime})\in{v_{l}}}\!\!\!\!c^{m}.$ (4)
Intuitively, Eq. (4) quantifies the tradeoff between predictive accuracy and the acquisition cost for candidate $v_{l}$, and we have access to $x_{v_{l}}$ since $x_{v_{l}}$ is from the training set. We encode each $v_{l}$ (including termination) by a binary vector $q_{l}\in\{0,1\}^{ML+1}$, where the first $ML$ entries are indexed by feature-time pairs in $v_{l}$ (e.g., $[q_{l}]_{(m,t^{\prime})}=\mathbf{1}_{(m,t^{\prime})\in v_{l}}$), and the last entry marks termination (e.g., $[q_{l}]_{ML+1}=\mathbf{1}_{v_{l}=\emptyset}$). We then define the probability that represents the distribution over the feature-time pair for each future candidate $v_{l}$ as $\bar{q_{l}}=\frac{q_{l}}{\|q_{l}\|_{1}}$, which is then used to define the future-candidate distribution below.
We define the future-candidate distribution $p(\cdot\mid x_{o})$ over feature-time pair (and termination) for the partial observed features $x_{o}$ as the weighted sum of the candidate distribution $\bar{q_{l}}$:
$p(\cdot\mid x_{o})=\sum_{l=1}^{\kappa}w_{l}\,\bar{q}_{l},$ (5)
where $w_{l}=\frac{\exp(-r_{l})}{\sum_{k=1}^{\kappa}\exp(-r_{k})}$. This probability vector $p(\cdot\mid x_{o})$ (Eq. (5)) indicates how likely each feature at each timestep (and the termination) is chosen, given a partially observed feature $x_{o}$. Note that applying Eq. (5) to specific $(x_{i},y_{i})\in\mathcal{D}$ having observed feature $o$ will yeild its future-candidate distribution $p(\cdot\mid x_{i,o})$.
Embedding Network. We train an embedding network $g_{\phi}:x_{i,o}\mapsto\mathbb{R}^{E}$ so that observations training $x_{i,o}$ and $x_{j,o}$, whose distributions $p(\cdot\mid x_{i,o})$ and $p(\cdot\mid x_{j,o})$ are close (by Jensen-Shannon (JS) divergence) lie nearby in embedding space $\mathbb{R}^{E}$, and those with dissimilar distributions lie at least a $\gamma>0$ margin apart. To quantify similarity, we use the exponential of the negative JS divergence between these two distributions:
$\texttt{sim}_{ij}=\exp\!\bigl{(}-\beta\,\mathrm{JS}(p(\cdot\mid x_{i,o}),\,p(\cdot\mid x_{j,o}))\bigr{)},$ (6)
where $\beta>0$ controls how strongly divergence affects similarity. Letting $\delta_{ij}=\bigl{\lVert}g_{\phi}\bigl{(}x_{i,o}\bigr{)}-g_{\phi}\bigl{(}x_{j,o}\bigr{)}\bigr{\rVert}^{2}$ be the distance between embeddings of two partial observations $\bigl{(}x_{i,o},x_{j,o}\bigr{)}$, we minimize a contrastive-type loss (Chopra et al., 2005):
$\mathcal{L}_{\mathrm{emb}}=\frac{1}{2}\Bigl{(}\texttt{sim}_{ij}\cdot\delta_{ij}+(1-\texttt{sim}_{ij})\cdot\max\bigl{(}0,\,\gamma-\delta_{ij}\bigr{)}^{2}\Bigr{)}.$ (7)
Intuitively, minimizing $\mathcal{L}_{\mathrm{emb}}$ (Eq. (7)) encourages pairs of samples whose future distributions are similar (low JS divergence) to have low Euclidean distance in the embedding space, while pairs with dissimilar distributions (high JS divergence) are pushed apart. Once $g_{\phi}$ has been trained, we retrieve the $K$-nearest neighbors (Eq. (3)) of any new observation $x_{o}$ by ranking all $(x_{i},y_{i})\in\mathcal{D}$ according to the embedding space distance as $d(x_{o},x_{i,o})=\|x_{o}-x_{i,o}\|_{2}$.
3.4 NOCTA-P: Parametric Method
Above, we developed a nonparametric, neighbor-based approach to estimating our generalized NOCTA cost objective Eq. (1). It is also natural to consider a direct parametric approach, which estimates the cost objective using a network. We expound on our parametric approach, coined NOCTA-P, below.
Estimating $u\bigl{(}{x}_{o},o\bigr{)}$. In order to estimate $u(x_{o},o)$ in Eq. (1) parametrically, we attempt to calculate the utility for all potential subsets $v\subseteq\mathcal{V}$:
$I_{v}(x_{o},o)=\ell(x_{o}\cup x_{v},y,t)+\alpha\!\!\!\!\sum_{(m,t^{\prime})\in v}\!\!\!\!c^{m}.$ (8)
Since we have access to all data during training, this value can be directly calculated for all subsets $v$ for training instances $(x_{i},y_{i})$. However, this is not the case during evaluation when $x_{v}$ becomes unknown. Therefore, we estimate $I_{v}(x_{o},o)$ through a value network, which we describe below.
Value Network. The value network $f_{\theta}$ aims to predict the utility $I_{v}(x_{o},o)$ for all potential subsets. Due to the inaccessibility of $x_{v}$ during evaluation, we use the candidate masks $v$ in addition to all the observed features $x_{o}$ and the current time point $t$ for utility prediction, resulting in an output value of $f_{\theta}(x_{o},v,t)$. Since this output aims to directly estimate the utility of $I_{v}(x_{o},o)$, we optimize through the mean squared error as:
$\mathcal{L}_{\mathrm{value}}=\mathrm{MSE}\left(f_{\theta}(x_{o},v,t),I_{v}(x_{o},o)\right)\,.$ (9)
To avoid distribution shift during evaluation, we match the distribution of observed subsets $o$ during training to what shall be encountered at inference time. That is, at each acquisition step for both training and inference, we choose the next subset based on $u(x_{o},o)=\arg\min_{v\subseteq V}f_{\theta}(x_{o},v,t)$.
During training, we jointly optimize the prediction network $\hat{y}$ and the value network $f_{\theta}$ on these potential feature subsets, resulting in the final loss function as:
$\mathcal{L}_{\mathrm{NOCTA\text{-}P}}=\mathcal{L}_{\mathrm{pred}}+\lambda\cdot\mathcal{L}_{\mathrm{value}}\,,$ (10)
where $\lambda$ balances between the prediction network loss $\mathcal{L}_{\mathrm{pred}}$ and the value network loss $\mathcal{L}_{\mathrm{value}}$. We describe our prediction network in Sec. 4.3.
4 Experiments
4.1 Datasets
Synthetic. Inspired by the dataset introduced by (Kossen et al., 2022), we construct a synthetic dataset containing $N=8,000$ samples, and each sample includes $L=10$ time points. Each instance contains a sequence of digit feature that is summed at locations where counter feature indicates (with a zero value). That is, the output labels are comprised of the cumulative sum of the counter-indicated digits. We expound below.
The digit feature consists of $L$ randomly generated integers uniformly selected from $\{0,1,2\}$, e.g., $1010220110$. The counter feature is created by concatenating countdown sequences that begin with randomly chosen starting values from $\{0,1,2\}$ and truncating to maintain a fixed length of $L$. For example, selecting starting values $2$, $1$, $2$, and $1$ generates the sequences $210$, $10$, $210$, $10$, and the final counter feature is their concatenation $\underline{2}10\underline{1}0\underline{2}10\underline{1}0$ (of length $L$). Unlike (Kossen et al., 2022), we extend the labels of the dataset to provide predictions at every time point. Specifically, the label $y_{t}$ is defined as the cumulative sum of the digit values at all previous timesteps where the corresponding counter value is 00, e.g., $0011333444$ in the case of ${\mathrm{counter}\atop\mathrm{digit}}$ ${2101021010\atop 1010220110}$. The policy should begin by acquiring the first counter feature, then acquiring the digit at the anticipated 00 counter location, then the next counter, and so forth. We split the dataset into train/val/test sets with a ratio of $70/15/15$.
ADNI. The Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset [^1] https://adni.loni.usc.edu (Petersen et al., 2010) is a longitudinal, multi-center, observational study. Patients are categorized into cognitively unimpaired, mild cognitive impairment, and Alzheimer’s Disease (AD) patients. We follow (Qin et al., 2024) and use a total of $1,002$ patients with four biomarkers extracted from PET (FDG and AV45) and MRI (Hippocampus and Entorhinal) for a total of $L=12$ time points. We split the dataset train/val/test with a ratio of $64/16/20$. Each patient may have a different pattern of missing features or labels, and we only report the results based on those that have a diagnosis label.
OAI. The Osteoarthritis Initiative (OAI) dataset [^2] https://nda.nih.gov/oai/ includes data from $N=4,796$ patients with separate evaluations for the left and right knees. Each patient is monitored longitudinally for up to $96$ months with varying patterns of missing data and labels due to the nature of the study design and patient participation. We use the tabular data selected from (Chen et al., 2024) as well as joint space width, which contains a total of $27$ features per time point. We evaluate our method for the prediction of two different clinical scores: the Kellgren-Lawrence grade (KLG) (Kellgren et al., 1957) (range from $0\sim 4$) and the Western Ontario and McMaster Universities (WOMAC) pain score (McConnell et al., 2001) (range from $0\sim 20$). Following previous works (Chen et al., 2024; Nguyen et al., 2024), we merge KLG $=0$ and $1$, and define WOMAC $<5$ as no pain and the $\geq 5$ as pain. We consider a total of $L=7$ visits and split the dataset into train/val/test sets by patient with a ratio of $50/12.5/37.5$. Appendix B.1 provides further details.
4.2 Baselines
We expand on comparisons performed in recent work by Qin et al. (2024) in longitudinal AFA settings by additionally comparing to two other state-of-the-art (SOTA) AFA methods, DIME (Gadgil et al., 2024) and DiFA (Ghosh and Lan, 2023), in addition to the RL method considered by Qin et al. (2024).
In total, we consider the following baselines: 1) ASAC (Yoon et al., 2019): uses the actor-critic method to jointly train a feature selector and a predictor network; 2) RAS (Qin et al., 2024): uses RL to decide when and which features to capture for longitudinal data using a continuous interval; 3) AS (Qin et al., 2024): RAS with a constant acquisition interval; 4) DIME (Gadgil et al., 2024): uses a greedy method to sequentially select the most informative features; 5) DiFA (Ghosh and Lan, 2023): uses a Gumbel-softmax-based differentiable policy for AFA.
To adapt DIME and DiFA for longitudinal data, we extend both methods by limiting the next capturing feature to either the current or future time points. Note that for DIME and DiFA, features are acquired one at a time, allowing repeated selection within the same time step $t$, while NOCTA and other baselines simultaneously select all desired features at $t$ before moving to the next time steps. This gives both DiFA and DIME an advantage by allowing them to decide whether they want to continue acquiring features at the current visit or advance to the next time step.
4.3 Implementation
[…]
$\mathcal{L}_{\mathrm{pred}}=\frac{1}{L}\sum_{k=1}^{L}{\mathrm{CE}\bigl{(}\hat{y}_{k}({x}_{o_{1:t}}),y_{k}\bigr{)}}.$ (11)
For predictions at time $t$ and for $k>t$, the network must predict label $y_{k}$ without access to any additional features beyond those observed by time $t$. We pretrain the prediction network with random dropout of the input features.
NOCTA-NP Implementation. For NOCTA-NP, we first pre-train the MLP predictor $\hat{y}_{k}$, followed by training the MLP-based embedding network $g_{\phi}$. Both networks are optimized with Adam (Kingma, 2014) with an initial learning rate of $31e-$ 3 for 20 epochs. For all the experiments, we train the embedding network with $\beta=1$, $\gamma=1$, and $\kappa=5$. Moreover, we use $K=5$ nearest neighbors and randomly sample $1000$ potential subsets.
NOCTA-P Implementation. We first pre-train the MLP prediction network $\hat{y}_{k}$, followed by jointly fine-tuning the prediction network and the MLP-based value network $f_{\theta}$. Both steps use an initial learning rate of $31e-$ 3 and Adam (Kingma, 2014) as the optimizer. We set $\lambda=1$ and use a random sample of $1000$ potential subsets.
Results are reported as the mean over five independent runs (standard errors for each run are provided in Appendix C). Further details on the experimental setup can be found in Appendix B. Moreover, Appendix A.5 shows a hyperparameter sensitivity analysis, demonstrating that our method remains robust across different settings and can be applied without extensive fine-tuning. Upon publication, we will make our code publicly available.
4.4 Performance-Cost Results
Figure 3: Performance/cost of models across various average acquisition costs (budgets). Following Kossen et al. (2022), we show accuracy rather than average precision for the synthetic dataset. Our NOCTA models show the highest accuracies for a given cost.
We show the result across different datasets in Fig. 3. For each figure, the x-axis represents different average acquisition costs (across instances), and the y-axis represents the performance. As the budget increases, more features can be acquired, and prediction performance typically improves. We use different hyperparameters to obtain prediction results under different average acquisition costs. Refer to Appendix B for more details on the trade-off hyperparameters used in our experiments.
Synthetic Dataset. Following (Kossen et al., 2022), we use accuracy to measure the performance on the synthetic dataset. NOCTA-NP outperforms the other baselines, followed by NOCTA-P, indicating that the baseline methods fail to model the conditional dependency between the features.
ADNI Dataset. Following (Qin et al., 2024), we show the average precision (AP) result for different costs. We can see that both the parametric and non-parametric versions of NOCTA consistently outperform all other baseline methods while using lower overall acquisition cost.
OAI Dataset. Similar to the ADNI results, we show the AP result on KLG and WOMAC prediction in Fig. 3. Our proposed NOCTA methods perform well and outperform the baselines. RL-based approaches (ASAC, AS, and RAS) did not perform as well as the greedy approach (DIME), indicating that RL-based frameworks may face challenges in handling complex patterns, i.e., more and noisier features, of longitudinal feature acquisition.
Overall, NOCTA-NP achieves slightly higher accuracy across benchmarks, whereas NOCTA-P is more computationally efficient. For additional results reporting AUC ROC and detailed time complexity and runtime analysis, please refer to Appendix C and Appendix A.3, respectively.
Table 1: Result comparisons of using the feature values, $\hat{y}$ prediction embedding (last hidden layer), and learned representation $g_{\phi}(\cdot)$ to compute nearest neighbors on datasets.
Method ADNI WOMAC KLG
AP ROC cost AP ROC cost AP ROC cost
Feature Values $0.545\pm 0.004$ $0.727\pm 0.002$ $7.052\pm 0.069$ $0.315\pm 0.003$ $0.617\pm 0.004$ $9.014\pm 2.863$ $0.629\pm 0.003$ $0.821\pm 0.002$ $11.024\pm 0.011$
$0.550\pm 0.007$ $0.732\pm 0.008$ $10.297\pm 1.168$ $0.331\pm 0.013$ $0.633\pm 0.013$ $17.574\pm 0.070$ $\mathbf{0.651}\pm\mathbf{0.010}$ $0.826\pm 0.021$ $17.741\pm 0.033$
Prediction Embedding $0.567\pm 0.005$ $0.744\pm 0.004$ $6.218\pm 0.038$ $0.290\pm 0.005$ $0.618\pm 0.002$ $6.998\pm 0.023$ $0.639\pm 0.004$ $\mathbf{0.825}\pm\mathbf{0.002}$ $13.706\pm 0.055$
$0.571\pm 0.021$ $0.744\pm 0.013$ $7.866\pm 0.068$ $0.290\pm 0.004$ $0.618\pm 0.002$ $19.681\pm 0.081$ $0.648\pm 0.001$ $\mathbf{0.829}\pm\mathbf{0.001}$ $24.705\pm 0.043$
Embedding Network $\mathbf{0.578}\pm\mathbf{0.004}$ $\mathbf{0.760}\pm\mathbf{0.003}$ $4.434\pm 0.042$ $\mathbf{0.328}\pm\mathbf{0.004}$ $\mathbf{0.646}\pm\mathbf{0.003}$ $8.729\pm 0.011$ $\mathbf{0.644}\pm\mathbf{0.003}$ $0.822\pm 0.001$ $10.687\pm 0.018$
$\mathbf{0.584}\pm\mathbf{0.013}$ $\mathbf{0.760}\pm\mathbf{0.006}$ $7.727\pm 0.094$ $\mathbf{0.339}\pm\mathbf{0.008}$ $\mathbf{0.647}\pm\mathbf{0.003}$ $20.921\pm 0.058$ $0.649\pm 0.002$ $0.822\pm 0.002$ $16.192\pm 0.062$
4.5 Ablation of NOCTA-NP Neighbor Distance
We evaluate the effectiveness of the learned representation during the nearest neighbors process. That is, we compare defining neighbors using the learned representation $g_{\phi}$ versus using the native feature values and using the last hidden layer from the predictor $\hat{y}$. For any two samples $x_{i}$ and $x_{j}$ having the same observed feature $o$, we define feature-based distance as: $d(x_{i,o},x_{j,o})=\|x_{i,o}-x_{j,o}\|_{2}$. Moreover, we define the prediction embedding using the last hidden layer of the predictor $\hat{y}$, denoted as $h(\cdot)$. Thus, the distance for the prediction embedding is computed as $d(x_{i,o},x_{j,o})=\|h(x_{i,o})-h(x_{j,o})\|_{2}$. Tab. 1 shows results under two cost regimes, low and high. Our non-parametric approach is robust to various choices of metrics and performs relatively well throughout. Moreover, the learned embedding $g_{\phi}(\cdot)$ outperforms both the feature values and the prediction embedding $h(\cdot)$ on both ADNI and WOMAC datasets, and it shows comparable results on the KLG dataset. Notably, on ADNI, the low-cost result for $g_{\phi}(\cdot)$ (cost $=4.343$) already outperforms the high-cost performance of both alternatives (cost $=7.866$ and $10.297$). This indicates the embedding network is able to learn informative representations that can be used for real-world AFA tasks.
5 Conclusion
In this work, we propose NOCTA, a non-greedy longitudinal active feature acquisition framework designed to directly balance acquisition cost with prediction accuracy. We introduce a cohesive estimation target for our NOCTA framework with two complementary solution strategies. Results on both synthetic and real-world medical datasets show that NOCTA outperforms other baseline methods while achieving lower overall acquisition cost. While our non-parametric approach provides consistently strong performance, the parametric approach provides comparable performance with more lightweight (non-neighbor based) inference. It is worth mentioning that either NOCTA-P or NOCTA-NP can be easily validated against held-out data or ensembles if one seeks a single policy.
6 Acknowledgment
This research was, in part, funded by the National Institutes of Health (NIH) under awards 1OT2OD038045-01, NIAMS 1R01AR082684, 1R01AA02687901A1, and 1OT2OD032581-02-321, and was partially funded by the National Science Foundation (NSF) under grants IIS2133595 and DMS2324394. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing official policies, either expressed or implied, of the NIH and NSF.
Appendix A Additional Discussion
A.1 Assumptions & Limitations
Like traditional AFA methods and RL-based baselines, NOCTA assumes that the problem has a finite horizon with a limited number of time steps. Additionally, we acknowledge that, similar to other baselines presented in this paper, the performance of NOCTA is sensitive to the data quality. In particular, poor label annotation in real-world clinical datasets can degrade the accuracy. Moreover, when the data is extremely sparse and noisy, methods may struggle to generate reasonable feature selection. However, extending to an infinite-horizon setting and mitigating the sensitivity to the data quality are beyond the scope of this paper. We will leave these challenges for future work.
A.2 Broader Impact
In this paper, we introduce a non-greedy longitudinal active feature acquisition method, NOCTA, which could aid the decision-making process in real-world applications such as healthcare by selectively acquiring essential examinations; thus minimizing patient burden (e.g., time and costs) through personalized and timely intervention, as in any ML system, using NOCTA should involve thorough/robust evaluations and supervision by medical professionals to prevent any potential negative impact.
Broader Application. We acknowledge that our experiments focus on the healthcare section, an important application area; however, our method is not specific to any domain. Similar resource-constrained scenarios also exist in other domains, such as in energy-aware audio recognition [Monjur et al., 2023] or IoT devices [Dunkels et al., 2011]. In fact, our AFA formulation has applications across a myriad of domains, including robotics (acquisition of sensors with resource constraints), education (acquisition of different exam types without over-burdening students), and beyond.
A.3 Time Complexity
[…]
A.4 Proof of Theorem 1
Theorem 1. The value of the optimal longitudinal AFA MDP policy is lower bounded by NOCTA.
Problem Setup. For any observations $(x_{o},o,t)$, let the set of all feature-time pairs that are available be:
$\mathcal{V}_{t}=\{(m,t^{\prime})\mid m\in\{1,\ldots,M\},\,t^{\prime}\in\{t+1,\ldots,L\}\}.$
To follow the convention of reinforcement learning literature, we redefine the objective for NOCTA (Eq. 1) as maximization instead of minimization without loss of generality. That is, for any partial observation $(x_{o},t)$, we define $\texttt{NOCTA}^{s}_{t}(x_{o})$ as:
$\texttt{NOCTA}_{t}^{s}(x_{o}):=\max_{\begin{subarray}{c}v\subseteq\mathcal{V}_{t}\\ \text{s.t }\mid\text{rounds}(v)\mid\leq s\end{subarray}}\Bigl{[}-\mathbb{E}_{y_{t+1:L},x_{v}\mid{x}_{o}}[\ell(x_{o}\cup x_{v},y,t)]-\alpha\!\!\!\!\sum_{(m,t^{\prime})\in v}\!\!\!\!c^{m}\Bigr{]},$
where $s\in\mathbb{N}$ represents the maximum number of possible future acquisition rounds, and let $\text{rounds}(v):=\{t^{\prime}\mid\exists m\in\{1,\ldots,M\}:(m,t^{\prime})\in v\}$. We then define
$V_{t}^{0}(x_{o}):=-\mathbb{E}_{y_{t+1:L}\mid x_{o}}[\ell(x_{o},y,t)]$
and
$V_{t}^{s}(x_{o})$ $:=\max\Bigl{\{}V_{t}^{0}(x_{o}),\max_{\begin{subarray}{c}t^{\prime}\in\{t+1,...,L\}\\ a_{t^{\prime}}\in\{0,1\}^{M}\end{subarray}}\Bigl{[}-\alpha\sum_{m}a_{t^{\prime},m}c^{m}+\mathbb{E}_{x_{a_{t^{\prime}}}\mid x_{o}}[V^{s-1}_{t^{\prime}}(x_{o\cup a_{t^{\prime}}})]\Bigr{]}\Bigr{\}}.$ (12)
We see that $V^{0}_{t}$ is the expected negative cumulative loss when we immediately terminate at $t$, and we predict the future labels using only what we have acquired so far $x_{o}$. Thus, $V^{s}_{t}$ is the value for the optimal longitudinal AFA policy, such that it either (i) terminates, returning $V_{t}^{0}$ or (ii) executes another acquisition at the future time step $t^{\prime}\in(t+1,...,L)$ and acquires the feature $a_{t^{\prime}}\in\{0,1\}^{M}$.
Claim. For every $(x_{o},t)$ with $s\in\mathbb{N}\geq 0$, we have $V_{t}^{s}(x_{o})\geq\texttt{NOCTA}^{s}_{t}(x_{o})$.
Proof.
Base case $s=0$. Since there is no further action, both the MDP agent and NOCTA terminate:
$V_{t}^{0}(x_{o})=-\mathbb{E}_{y_{t+1:L}\mid x_{o}}[\ell(x_{o},y,t)]=\texttt{NOCTA}^{0}_{t}(x_{o}).$
Inductive hypothesis. Assume for $s-1\geq 0$, we have:
$V_{t^{\prime}}^{s-1}(x)\geq\texttt{NOCTA}^{s-1}_{t^{\prime}}(x)\quad\forall(x,t^{\prime}).$ (IH)
Inductive step. We fix $(x_{o},t)$ and see that:
$V_{t}^{s}(x_{o})$
$=\max\Bigl{\{}V_{t}^{0}(x_{o}),\max_{\begin{subarray}{c}t^{\prime}\in\{t+1,...,L\}\\ a_{t^{\prime}}\in\{0,1\}^{M}\end{subarray}}\Bigl{[}-\alpha\sum_{m}a_{t^{\prime},m}c^{m}+\mathbb{E}_{x_{a_{t}^{\prime}}\mid x_{o}}[V^{s-1}_{t^{\prime}}(x_{o\cup a_{t^{\prime}}})]\Bigr{]}\Bigr{\}}$
$\!\!\overset{(\text{IH})}{\geq}\max\Bigl{\{}V_{t}^{0}(x_{o}),\max_{\begin{subarray}{c}t^{\prime}\in\{t+1,...,L\}\\ a_{t^{\prime}}\in\{0,1\}^{M}\end{subarray}}\Bigl{[}-\alpha\sum_{m}a_{t^{\prime},m}c^{m}+\mathbb{E}_{x_{a_{t}^{\prime}}\mid x_{o}}[\texttt{NOCTA}^{s-1}_{t^{\prime}}(x_{o\cup a_{t^{\prime}}})]\Bigr{]}\Bigr{\}}$
$=\max\Bigl{\{}V_{t}^{0}(x_{o}),$
$\max_{\begin{subarray}{c}t^{\prime}\in\{t+1,...,L\}\\ a_{t^{\prime}}\in\{0,1\}^{M}\end{subarray}}\Bigl{[}-\alpha\sum_{m}a_{t^{\prime},m}c^{m}+\mathbb{E}_{x_{a_{t^{\prime}}}\mid x_{o}}\Bigl{[}\!\!\!\!\max_{\begin{subarray}{c}v^{\prime}\subseteq\mathcal{V}_{t^{\prime}}\\ \text{s.t }\mid\text{rounds}(v^{\prime})\mid\leq s-1\end{subarray}}\!\!\!\!\!\!\!\!\!\!\!\Bigl{(}-\mathbb{E}_{y_{t^{\prime}+1:L},x_{v^{\prime}}\mid{x}_{o},x_{a_{t^{\prime}}}}[\ell(x_{o}\cup x_{a_{t^{\prime}}}\cup x_{v^{\prime}},y,t^{\prime})]-\alpha\!\!\!\!\!\!\sum_{(m,t^{\prime\prime})\in v^{\prime}}\!\!\!\!\!\!c^{m}\Bigr{)}\Bigr{]}\Bigr{]}\Bigr{\}}$
$\overset{*}{\geq}\max\Bigl{\{}V_{t}^{0}(x_{o}),$
$\max_{\begin{subarray}{c}t^{\prime}\in\{t+1,...,L\}\\ a_{t^{\prime}}\in\{0,1\}^{M}\end{subarray}}\max_{\begin{subarray}{c}v^{\prime}\subseteq\mathcal{V}_{t^{\prime}}\\ \text{s.t }\mid\text{rounds}(v^{\prime})\mid\leq s-1\end{subarray}}\left[-\mathbb{E}_{y_{t^{\prime}+1:L},x_{v^{\prime}},x_{a_{t^{\prime}}}\mid{x}_{o}}[\ell(x_{o}\cup x_{a_{t^{\prime}}}\cup x_{v^{\prime}},y,t^{\prime})]-\alpha\Bigl{(}\sum_{(m,t^{\prime\prime})\in v^{\prime}}\!\!\!\!c^{m}+\sum_{m}a_{t^{\prime},m}c^{m}\Bigr{)}\Bigr{\}}\right]$
$=\max\Bigl{\{}V_{t}^{0}(x_{o}),\max_{\begin{subarray}{c}v\in\mathcal{V}_{t}\\ \text{s.t }\mid\text{rounds}(v)\leq s\mid\end{subarray}}[-\mathbb{E}_{y_{t+1:L},x_{v}\mid{x}_{o}}[\ell(x_{o}\cup x_{v},y,t)]-\alpha\!\!\!\!\sum_{(m,t^{\prime\prime})\in v}\!\!\!\!c^{m}]\Bigr{\}}$
$=\texttt{NOCTA}^{s}_{t}(x_{o}).$
where * follows from:
$\begin{array}[]{r}\forall v^{\prime}\in\Omega,x_{a_{t^{\prime}}},\max\limits_{v\in\Omega}\mathcal{L}\left(x_{a_{t^{\prime}}},v\right)\geq\mathcal{L}\left(x_{a_{t^{\prime}}},v^{\prime}\right)\quad\Longrightarrow\\[8.00003pt] \forall v^{\prime}\in\Omega,\mathbb{E}_{x_{a_{t^{\prime}}}\mid x_{o}}\left[\max\limits_{v\in\Omega}\mathcal{L}\left(x_{a_{t^{\prime}}},v\right)\right]\geq\mathbb{E}_{x_{a_{t^{\prime}}}\mid x_{o}}\left[\mathcal{L}\left(x_{a_{t^{\prime}}},v^{\prime}\right)\right]\Longrightarrow\\[8.00003pt] \mathbb{E}_{x_{a_{t^{\prime}}}\mid x_{o}}\left[\max\limits_{v\in\Omega}\mathcal{L}\left(x_{a_{t^{\prime}}},v\right)\right]\geq\max\limits_{v\in\Omega}\mathbb{E}_{x_{a_{t^{\prime}}}\mid x_{o}}\left[\mathcal{L}\left(x_{a_{t^{\prime}}},v\right)\right]\end{array}$
for $\Omega:=\Bigl{\{}v\subseteq\mathcal{V}_{t^{\prime}}\;\bigm{|}\;|\text{rounds}(v)|\leq s-1\Bigr{\}}$ and
$\mathcal{L}(x_{a_{t^{\prime}}},v):=-\mathbb{E}_{y_{t+1:L},x_{v^{\prime}}\mid{x}_{o},x_{a_{t^{\prime}}}}[\ell(x_{o}\cup x_{a_{t^{\prime}}}\cup x_{v^{\prime}},y,t)]-\alpha\left(\sum_{(m,t^{\prime\prime})\in v^{\prime}}\!\!\!\!c^{m}+\sum_{m}a_{t^{\prime},m}c^{m}\right).$
Note that one may recover the non-cardinality constrained problem by considering large enough $s$, yielding a lower bound on the longitudinal AFA MDP.
A.5 Sensitivity Analysis
In this section, we provide further explanations for hyperparameters and how the hyperparameters affect performance. Overall, we find that our proposed method is robust to different settings, indicating that our method can be applied readily without excessive fine-tuning.
[…]
$\kappa$: this controls how many top candidate subsets are aggregated to form the future acquisition distribution. Please refer to Table 5 for different settings.
$K$: this controls how many neighbors in the learned embedding space are used to estimate future acquisition gains. Small values of $K$ might result in high variance from too few neighbors, while high $K$ may include dissimilar samples. Please refer to Table 4 for different settings.
$\beta$: this controls the sensitivity that translates Jessen-Shannon divergence into a similarity score. Larger $\beta$ will penalize divergences more strongly, while smaller $\beta$ will be more lenient. Please refer to Table 6 for different settings.
$\gamma$: this determines the margin in the contrastive loss, making sure that dissimilar embeddings are pushed far apart. If $\gamma$ is too small, the embedding network might not push dissimilar pairs far enough, and if $\gamma$ is too large, it might push moderately different samples by a large distance. Please refer to Table 7 for different settings.
[…]
$100$ $0.907\pm 0.001$ $13.252\pm 0.044$
$1000$ $0.919\pm 0.002$ $11.345\pm 0.012$
$5000$ $0.921\pm 0.001$ $12.046\pm 0.021$
$10000$ $0.921\pm 0.002$ $11.945\pm 0.012$
[…]
$100$ $0.731\pm 0.010$ $11.401\pm 0.040$
$1000$ $0.745\pm 0.006$ $11.624\pm 0.087$
$5000$ $0.749\pm 0.016$ $11.430\pm 0.115$
$10000$ $0.742\pm 0.017$ $11.419\pm 0.150$
Table 4: Ablation on the number of neighbors $K$ used in the synthetic dataset
$K$ Accuracy Cost
$5$ $0.888\pm 0.005$ $9.471\pm 0.033$
$10$ $0.874\pm 0.003$ $9.489\pm 0.006$
$25$ $0.889\pm 0.001$ $9.512\pm 0.006$
Table 5: Ablation on the number of top candidate subset $\kappa$ used in the synthetic dataset
$\kappa$ Accuracy Cost
$5$ $0.884\pm 0.005$ $9.471\pm 0.033$
$10$ $0.883\pm 0.001$ $9.618\pm 0.022$
$25$ $0.880\pm 0.002$ $9.564\pm 0.067$
Table 6: Ablation on the scaling in similarity function $\beta$ used in the synthetic dataset
$\beta$ Accuracy Cost
$0.5$ $0.879\pm 0.002$ $9.660\pm 0.054$
$1$ $0.884\pm 0.005$ $9.471\pm 0.033$
$2$ $0.890\pm 0.001$ $9.592\pm 0.022$
$5$ $0.888\pm 0.003$ $9.640\pm 0.048$
Table 7: Ablation on the margin in contrastive loss $\gamma$ used in the synthetic dataset
$\gamma$ Accuracy Cost
$0.5$ $0.882\pm 0.010$ $9.592\pm 0.035$
$1$ $0.884\pm 0.005$ $9.471\pm 0.033$
$2$ $0.879\pm 0.005$ $9.638\pm 0.033$
$5$ $0.880\pm 0.001$ $9.688\pm 0.017$
Appendix B Experimental Setup
B.1 Cost of Features
Synthetic data. Our synthetic dataset includes $2$ features, one being the digit and the other being the counter. Since both are generated through random selection, we assign an equal cost, $1$, to the two features. Upon publication, we will release the synthetic dataset.
ADNI. The ADNI dataset includes $4$ features per time point, with FDG and AV45 extracted from PET scan and Hippocampus and Entorhinal extracted from MRI scan. Since PET is a more expensive diagnosis, we follow [Qin et al., 2024] and assign cost $1$ to FDG and AV45, and cost $0.5$ to Hippocampus and Entorhinal. Access to ADNI data may be requested via https://adni.loni.usc.edu, and the Data Use Agreement is available at https://adni.loni.usc.edu/terms-of-use/.
OAI. We use a total of $27$ features for the OAI dataset, with $17$ clinical measurements, and $10$ joint space width (JSW) extracted from knee radiography. For the clinical measurements, we assign a low cost of $0.3$ to those that require minimum effort, e.g., age, sex, and race, and a slightly higher cost of $0.5$ to blood pressure and BMI calculation. For JSW, a cost of $1.0$ is assigned for the minimum JSW and $0.8$ for those measured at different positions. Access to the OAI data may be requested via https://nda.nih.gov/oai/.
B.2 Baseline Implementation
Follow  [Qin et al., 2024], we share the same neural CDE predictor  [Kidger et al., 2020] for ASAC, RAS, and AS. For ASAC, RAS, and AS, we use the implementation available at https://github.com/yvchao/cvar_sensing, which is under the BSD-3-Clause license. For DIME [Gadgil et al., 2024], we extend the method to longitudinal by building on top of the official implementation available at https://github.com/suinleelab/DIME. The DiFA codebase is not publicly released, and researchers may request it directly from the authors.
ASAC [Yoon et al., 2019]. We select the coefficient for the acquisition cost: 1) $\mu\in\{0.0015,0.002,0.003\}$ for the synthetic dataset, 2) $\mu\in\{0.005,0.02\}$ for ADNI, 3) $\mu\in\{0.0012,0.0016,0.003\}$ for the WOMAC of OAI dataset, and 4) $\mu\in\{0.00125,0.0015,0.00175,0.005\}$ for the KLG of OAI dataset.
RAS [Qin et al., 2024]. For the synthetic dataset, we set the allowed acquisition interval to $(\Delta_{\min}=0.2,\Delta_{\max}=1.0)$ for the synthetic dataset and $(\Delta_{\min}=0.5,\Delta_{\max}=4.5)$ for ADNI, WOMAC, and KLG.
Moreover, we use the coefficient for diagnostic error: 1) $\gamma\in\{5000,7000,9000\}$ for the synthetic dataset, 2) $\gamma\in\{50,75,100,175\}$ for ADNI dataset, 3) $\gamma\in\{2000,4500,9000\}$ for WOMAC of OAI, and 4) $\gamma\in\{1200,1700\}$ for KLG of OAI dataset. For all datasets, we use a tail-risk quantile of $0.1$, a penalty for invalid visits of $10$, and a discount factor (for tackling long trajectories) of $0.99$.
AS [Qin et al., 2024]. We set the minimum and maximum allowed acquisition intervals as in the RAS setting. In addition, we select fixed acquisition interval of: 1) $\tilde{\Delta}\in\{0.2,0.4\}$ for the synthetic dataset, 2) $\tilde{\Delta}\in\{0.5,1.0,1.5\}$ for ADNI dataset, 3) $\tilde{\Delta}\in\{0.5,1.5\}$ for WOMAC of OAI, and 4) $\tilde{\Delta}\in\{0.5,1.5\}$ for KLG of OAI dataset.
DIME [Gadgil et al., 2024]. For this baseline, we extend the original model to the longitudinal setting, while still keeping its greedy feature acquisition strategy. Both the prediction network and value network share the same architecture as our NOCTA-P model, and hence we use the same training strategies.
DiFA [Ghosh and Lan, 2023]. Following  [Ghosh and Lan, 2023], we use a variational autoencoder with arbitrary conditioning (VAEAC) probabilistic model [Ivanov et al., 2018] for imputation. When using the same architecture with NOCTA-P, DiFA performs poorly. Hence, we gave DiFA an advantage and used the recommended architecture from the DiFA paper. That is, both the prediction network and the feature policy model have two fully connected layers (hidden size 128) with skip connections, dropout regularizer, and LeakyReLU activation function.
B.3 NOCTA Implementation
Prediction Network. We share the same MLP architecture for both NOCTA-NP and NOCTA-P. Specifically, the network consists of two hidden layers with ReLU. Each hidden layer contains $10$ units for the synthetic and ADNI datasets, and $32$ units for the WOMAC and KLG of the OAI dataset, respectively. Moreover, we apply random dropout for the input during training the predictor.
NOCTA-NP. We select the accuracy-cost trade-off hyperparameter: 1) $\alpha\in\{0.002,0.006,0.02,0.03\}$ for the synthetic dataset, 2) $\alpha\in\{0.003,0.005,0.01\}$ for ADNI dataset, 3) $\alpha\in\{0.00075,0.005,0.01\}$ for WOMAC of OAI dataset, and 4) $\alpha\in\{0.0025,0.001,0.006\}$ for KLG of OAI dataset.
For the embedding network, we use four hidden layers with ReLU. Each hidden layer contains $32$ units for the synthetic and ADNI datasets, and $64$ units for the WOMAC and KLG of the OAI dataset, respectively.
NOCTA-P. For the value network, we use MLP architecture with two hidden layers and ReLU as the activation function. We use the same hidden size as the prediction network, but the number of output is set to be $1$ for estimating the utility of the potential subset. We set the accuracy-cost trade-off hyperparameter $\alpha$ to $55e-$ 5 for WOMAC and $45e-$ 4 to all other tasks.
To promote more stable training for Eq. (8) in NOCTA-P, we slightly modify the value network objective in two ways: 1) remove the acquisition cost and train the value network to estimate only the CE loss, i.e., $\bar{I}_{v}(x_{o},o)=I_{v}(x_{o},o)-\alpha\sum_{(m,t^{\prime})\in v}c^{m}$. Since the cost of each subset is fixed and can be directly added during subset selection, excluding it from the prediction target simplifies the learning process without affecting the final decision; 2) consider the relative prediction improvement over the current feature set: $\bar{\bar{I}}_{v}(x_{o},o)=\ell(x_{o},y,t)-\bar{I}_{v}(x_{o},o)$. This represents the potential gain from acquiring new features and implicitly normalizes the loss scale. Since $\ell(x_{o},y,t)$ is a constant, this will not change the selection (i.e., through $\arg\max$), and hence remains the same objective as Eq. (1).
B.4 Hardwares
We employed two different clusters for our experiments: (1) Intel Xeon CPU E5-2630 v4 and NVIDIA GeForce GTX 1080 for experiments and (2) Intel Xeon Silver 4114 CPU and NVIDIA GeForce GTX 2080. We used cluster (1) for RAS, AS, ASAC, DiFA, NOCTA-NP, and cluster (2) for DIME and NOCTA-P.
Appendix C Additional Results
C.1 ROC AUC on Real-World Datasets
Following [Qin et al., 2024], Figure 4 shows additional ROC AUC results on real-world datasets across varying average acquisition budgets. Our method consistently outperforms all baselines while using lower cost.
Figure 4: Additional results on performance/cost of models, measured by ROC AUC across varying average acquisition budgets on real-world datasets.
C.2 Full Results with Mean and Standard Deviation
Tables 8, 9, 10, and 11 report the mean and standard deviation for each data point presented in Figures 3 and 4, computed over five independent runs.
Table 8: Full results for accuracy vs. average cost on Synthetic Dataset.
Method Synthetic
Accuracy Cost
ASAC $0.449\pm 0.078$ $5.803\pm 5.180$
$0.567\pm 0.161$ $9.929\pm 8.365$
$0.624\pm 0.136$ $11.780\pm 5.211$
RAS $0.316\pm 0.031$ $6.390\pm 5.316$
$0.315\pm 0.032$ $6.852\pm 5.129$
$0.336\pm 0.022$ $11.003\pm 2.638$
AS $0.320\pm 0.014$ $9.210\pm 2.344$
$0.335\pm 0.033$ $12.870\pm 3.762$
DIME $0.497\pm 0.078$ $4.000\pm 0.000$
$0.501\pm 0.080$ $4.662\pm 0.557$
$0.509\pm 0.084$ $5.375\pm 1.213$
DiFA $0.444\pm 0.015$ $6.466\pm 0.058$
$0.463\pm 0.017$ $8.249\pm 0.420$
$0.504\pm 0.010$ $10.138\pm 0.129$
NOCTA-NP $0.698\pm 0.008$ $5.028\pm 0.022$
$0.795\pm 0.002$ $6.504\pm 0.005$
$0.884\pm 0.005$ $9.471\pm 0.033$
$0.919\pm 0.002$ $11.345\pm 0.012$
NOCTA-P $0.512\pm 0.018$ $3.580\pm 0.292$
$0.589\pm 0.022$ $5.666\pm 0.221$
$0.630\pm 0.026$ $7.697\pm 0.174$
$0.707\pm 0.017$ $9.624\pm 0.164$
$0.745\pm 0.006$ $11.624\pm 0.087$
Table 9: Full results for AP and ROC vs. average cost on ADNI prediction.
Method ADNI
AP ROC Cost
ASAC $0.449\pm 0.041$ $0.657\pm 0.038$ $6.442\pm 6.093$
$0.497\pm 0.034$ $0.679\pm 0.015$ $12.242\pm 2.648$
RAS $0.538\pm 0.036$ $0.717\pm 0.024$ $4.087\pm 0.735$
$0.567\pm 0.015$ $0.735\pm 0.008$ $5.064\pm 0.324$
$0.567\pm 0.022$ $0.737\pm 0.012$ $6.560\pm 2.459$
$0.572\pm 0.012$ $0.739\pm 0.011$ $9.197\pm 1.438$
AS $0.514\pm 0.046$ $0.702\pm 0.038$ $5.133\pm 1.659$
$0.536\pm 0.021$ $0.721\pm 0.014$ $6.394\pm 2.477$
$0.546\pm 0.021$ $0.734\pm 0.013$ $12.054\pm 1.763$
DIME $0.519\pm 0.037$ $0.671\pm 0.047$ $1.801\pm 0.050$
$0.538\pm 0.045$ $0.695\pm 0.048$ $3.607\pm 0.084$
$0.544\pm 0.048$ $0.703\pm 0.049$ $4.587\pm 0.330$
DiFA $0.520\pm 0.027$ $0.717\pm 0.029$ $2.925\pm 0.127$
$0.529\pm 0.025$ $0.725\pm 0.028$ $4.180\pm 0.043$
$0.530\pm 0.027$ $0.727\pm 0.033$ $5.292\pm 0.107$
NOCTA-NP $0.578\pm 0.004$ $0.760\pm 0.003$ $4.434\pm 0.042$
$0.579\pm 0.009$ $0.760\pm 0.003$ $6.164\pm 0.068$
$0.584\pm 0.013$ $0.760\pm 0.006$ $7.727\pm 0.094$
NOCTA-P $0.548\pm 0.025$ $0.719\pm 0.025$ $1.851\pm 0.019$
$0.569\pm 0.023$ $0.730\pm 0.015$ $3.729\pm 0.079$
$0.576\pm 0.021$ $0.742\pm 0.012$ $6.743\pm 0.330$
Table 10: Full results for AP and ROC vs. average cost on WOMAC score prediction.
Method WOMAC
AP ROC Cost
ASAC $0.269\pm 0.032$ $0.586\pm 0.060$ $8.541\pm 2.136$
$0.293\pm 0.018$ $0.619\pm 0.028$ $21.423\pm 6.480$
$0.299\pm 0.025$ $0.626\pm 0.026$ $27.175\pm 6.933$
RAS $0.226\pm 0.008$ $0.534\pm 0.019$ $15.930\pm 3.475$
$0.231\pm 0.007$ $0.543\pm 0.009$ $18.810\pm 1.802$
$0.251\pm 0.019$ $0.565\pm 0.024$ $24.939\pm 3.796$
AS $0.225\pm 0.005$ $0.528\pm 0.008$ $9.160\pm 2.271$
$0.235\pm 0.011$ $0.536\pm 0.012$ $19.826\pm 10.326$
DIME $0.297\pm 0.025$ $0.598\pm 0.043$ $7.073\pm 0.093$
$0.297\pm 0.025$ $0.601\pm 0.043$ $9.036\pm 0.198$
$0.302\pm 0.028$ $0.603\pm 0.044$ $11.009\pm 0.310$
$0.305\pm 0.027$ $0.617\pm 0.037$ $12.696\pm 0.357$
$0.310\pm 0.028$ $0.621\pm 0.038$ $13.325\pm 0.643$
DiFA $0.199\pm 0.009$ $0.504\pm 0.004$ $9.532\pm 0.091$
$0.211\pm 0.005$ $0.502\pm 0.002$ $11.584\pm 0.630$
$0.217\pm 0.004$ $0.507\pm 0.008$ $14.936\pm 1.433$
NOCTA-NP $0.317\pm 0.003$ $0.636\pm 0.0012$ $7.063\pm 0.004$
$0.328\pm 0.004$ $0.646\pm 0.003$ $8.729\pm 0.011$
$0.339\pm 0.008$ $0.647\pm 0.003$ $20.921\pm 0.058$
NOCTA-P $0.311\pm 0.007$ $0.625\pm 0.006$ $11.862\pm 0.074$
$0.327\pm 0.020$ $0.641\pm 0.018$ $13.212\pm 0.180$
$0.335\pm 0.015$ $0.649\pm 0.015$ $18.178\pm 0.082$
$0.340\pm 0.013$ $0.658\pm 0.011$ $23.061\pm 0.059$
Table 11: Full results for AP and ROC vs. average cost on KLG prediction.
Method KLG
AP ROC Cost
ASAC $0.258\pm 0.021$ $0.513\pm 0.035$ $7.284\pm 4.694$
$0.266\pm 0.018$ $0.519\pm 0.019$ $12.756\pm 6.584$
$0.283\pm 0.009$ $0.546\pm 0.024$ $20.025\pm 6.504$
$0.304\pm 0.008$ $0.589\pm 0.015$ $24.361\pm 15.940$
RAS $0.290\pm 0.012$ $0.571\pm 0.023$ $20.253\pm 11.269$
$0.313\pm 0.049$ $0.593\pm 0.063$ $31.044\pm 8.560$
AS $0.277\pm 0.012$ $0.552\pm 0.021$ $12.394\pm 2.027$
$0.285\pm 0.022$ $0.555\pm 0.015$ $36.499\pm 8.124$
DIME $0.418\pm 0.016$ $0.672\pm 0.023$ $9.787\pm 0.040$
$0.438\pm 0.018$ $0.688\pm 0.028$ $12.281\pm 0.005$
$0.486\pm 0.026$ $0.734\pm 0.024$ $14.359\pm 0.148$
DiFA $0.310\pm 0.006$ $0.595\pm 0.019$ $9.611\pm 1.743$
$0.322\pm 0.011$ $0.598\pm 0.015$ $22.331\pm 3.465$
$0.336\pm 0.010$ $0.606\pm 0.013$ $27.615\pm 6.411$
NOCTA-NP $0.644\pm 0.003$ $0.822\pm 0.001$ $10.687\pm 0.018$
$0.649\pm 0.002$ $0.822\pm 0.002$ $16.192\pm 0.062$
$0.663\pm 0.004$ $0.831\pm 0.001$ $29.546\pm 0.052$
NOCTA-P $0.641\pm 0.054$ $0.844\pm 0.020$ $13.144\pm 0.107$
$0.639\pm 0.056$ $0.845\pm 0.022$ $16.448\pm 0.228$
$0.655\pm 0.042$ $0.853\pm 0.017$ $21.776\pm 0.323$